OpenStudy (anonymous):

Solve. 3(x + 1) - 2x = -6.

OpenStudy (texaschic101):

3(x + 1)- 2x = - 6 (distribute the 3 through the parenthesis) 3x + 3 - 2x = - 6 (combine like terms) x + 3 = - 6 (subtract 3 from both sides) x + 3 - 3 = - 6 -3 (combine like terms) x = - 9 check... 3(x + 1) - 2x = - 6 3(-9 + 1) - 2(-9) = - 6 3(-8) + 18 = - 6 -24 + 18 = - 6 -6 = -6 (correct) ANSWER : x = - 9